OpenStudy (anonymous):

Find f(g(x)) how do you find g(x) = 2x^2 + 4 and f(x) = x-2/4 ?

OpenStudy (anonymous):

replace the x in f(x) with 2x^2 + 4

OpenStudy (anonymous):

Then simplify

OpenStudy (anonymous):

okay so do 2x^2+4-2/4?

OpenStudy (anonymous):

\[f(x) = x - \frac{2}{4}\ = x- \frac{1}{2}\] \[f((g(x)) = ( 2x^2 + 4 ) - \frac{1}{2} \] Unless you didn't write f(x) correctly and meant... \[f(x) = \frac{(x-2)}{4}\] (Note: if that's what you meant to write, then you HAVE TO make use of parenthesis or order of operations takes precedence and division goes first before subtraction)

OpenStudy (anonymous):

hahah @agentx5 you were just helping me out.. thats funny! @noniebug is my sister :)

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

hehe i need help. im not a smart girl :(

OpenStudy (anonymous):

no it has no "()" around it..

OpenStudy (anonymous):

its just x-2/4..

OpenStudy (anonymous):

I hope this makes sense then, missed parenthesis causing order of operations not to be what people intended is the #1 syntax error people probably make on this section of the website. But it's very important if people are trying to answer, because it can completely change what was meant, as you can see from above: 2/4 reduces to a half. Think of it this way: g squares the input, then doubles it, then adds four f subtracts whatever the input is by one half Best answer? :D

OpenStudy (anonymous):

oh my gosh. like i said im not a smart girl so i hardly got any of that :/ haaa.

OpenStudy (anonymous):

What part are you having trouble understanding?

OpenStudy (anonymous):

the whole thing..

OpenStudy (anonymous):

i already forgot what question i askedd...